OpenStudy (anonymous):

Select the best possible first step to solving the following system by first eliminating the x variable. 2x + 5y = 1 x + 6y = –10

OpenStudy (anonymous):

multiply the bottom one by 2, then subtract the bottom from top or vice versa.

hero (hero):

set x = x x = (1-5y)/2 x = (10 + 6y)

OpenStudy (anonymous):

twoof and Hero are both correct.

OpenStudy (anonymous):

2x+5y = 1 x + 6y = -10 I would multiply the second equation by -2, and then add them together, using the elimination method. 2x + 5y = 1 -2x - 12y = 20 Which simplifies to: -7y = 21 y = -3 Now substitute in the y value into either equation: 2x + 5(-3) = 1 2x - 15 =1 2x = 16 x = 8 Now in the second equation. x + 6y = -10 x + 6(-3) = -10 x + -18 = -10 x = 8 Hope this helps. http://www.tutorsean.net